ANOVA Source of Variation Factor A Factor B Interaction Error p-value MS 9,492.72 11,533.52 19,995.89 7,629.83 df 28,478.16 23,067.03 119,975.34 274,673.79 36 Total 446,194.32 47 (a) What kind of ANOVA is this? One-factor ANOVA O Two-factor ANOVA with replication O Two-factor ANOVA without replication (b) Calculate each F test statistic and the p-value for each Ftest using Excers function -F.DISTRT(FDF1,DF2) (Round your Fcalc values to 3 decimal places and p-values to 4 decimal places.) Source of Variation Factor A Factor B Interaction p-value (c-1) There is a significant effect due to Factor A. O False True (c-2) There is a significant effect due to Factor B 29 MacBook 80 F3 O00 F4 F1 F2 FS F6
Answer #1

your need of answer of part (b)

Source of Variation F p-value
Factor A 1.244 0.3081
Factor B 1.512 0.2342
Interaction 2.621 0.0327

F is the ratio of MS(effect) and MS(error)

for factor A the F=9492.72/7629.83=1.244 and its p-value=0.3081 ( using ms-excel=fdist(1.244,3,36)

for factor B the F=11533.52/7629.83=1.512 and its p-value=0.2342 ( using ms-excel=fdist(1.512,2,36)

for interaction the F=19995.89/7629.83=2.621 and its p-value=0.2342 ( using ms-excel=fdist(1.512,6,36)

part(a) and part(c) is true